Melanie recently hired a plumber to do some necessary work. On the final bill, Melanie was charged a total of $839. $210 was listed for parts and the rest for labor. If the hourly rate for labor was $34, how many hours of labor was needed to complete the job?
(A) First write an equation you can use to answer this question. Use
x as your variable and express any percents in decimal form in the equation.
The equation is_____

(B) Solve your equation in part (A) to find the number of labor hours needed to do the job.
Answer: The number of labor hours was___
.

Found 2 solutions by ankor@dixie-net.com, mananth:
Answer by ankor@dixie-net.com(22740)

On the final bill, Melanie was charged a total of $839.
$210 was listed for parts and the rest for labor.
If the hourly rate for labor was $34, how many hours of labor was needed to complete the job?
:
(A) First write an equation you can use to answer this question. Use
x as your variable and express any percents in decimal form in the equation.
The equation is:
34x + 210 = 839
:
(B) Solve your equation in part (A) to find the number of labor hours needed to do the job.
Answer: The number of labor hours was:
34X + 210 = 839
34X = 839 - 210
34X = 629
x = 629/34
x = 18.5 hrs of labor
.

Answer by mananth(16946)


Let x be the number of hours the plumber worked
The equation is 839= 210+34x
839 -210=34x
629 =34x
x=18.5
The number of labor hours was 18.5